OpenStudy (anonymous):

if the volume of the cylinder =141cm^3,find the volume of the cone ?

OpenStudy (campbell_st):

divide the volume of the cylinder by 3 this is because a cone and cylinder with the same radius and height have the following volume formula cylinder \[V = \pi r^2 h\] cone \[V = \frac{1}{3} \pi r^2h\] hope it helps

OpenStudy (anonymous):

@campbell_st can you please go through it with me step by step ?

OpenStudy (anonymous):

@campbell_st Wait, so the answer is 47?